OpenStudy (anonymous):

What is the sum of the geometric sequence 1, -4, 16, if there are 7 terms? (1 points) -5,461 5,461 -3,277 3,277

OpenStudy (campbell_st):

can you tell me what the common ratio r is...?

OpenStudy (anonymous):

Huh ?

OpenStudy (campbell_st):

well a geometric sequence needs a common ratio the next term = current term x the common ratio.... so what did you multiply the 1st term by to get the 2nd term... does it work when you multiply the 2nd term by the same value to get the 3rd term

OpenStudy (anonymous):

You multiply -4 To get the next term

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

great thats the common ratio r = -4 and the 1st term is a = 1 then you need to find the sum of 7 terms so n = 7 and use the formula \[S_{n} = \frac{1 \times (1 - r^n)}{1 - r} \] make the necessary substitutions and you'll get the sum of 7 terms

OpenStudy (anonymous):

What ?

OpenStudy (whpalmer4):

Like the gent said, plug in the values of \(r, n\) to find \(S_n\) which is the sum of the first \(n\) terms of the geometric sequence with \(r = -4\) and \(a=1\). \[S_7 = \frac{1*(1-(-4)^7)}{1-(-4)} = \] You can check your work by finding the first 7 terms and adding them up.